# Context Matters: Learning Generalizable Rewards via Calibrated Features

### v2

@@ D.5 Procedures
At the start of each query response round, participants were primed with a general preference, prompting them to focus on a specific contextual element to teach the effect of. Participants were then asked to write down their personal contextual preference (see prompt example for stove distance feature in Fig. 20). Again, while our method supports accounting for the contextual effects of numerous contextual elements, we used this prompting to steer participants to teach the contextual effects of a single element such that we could visualize changes in the affecting contextual element when they visualized the learned models. The query response GUI was then pulled up (Fig. 18), and participants were prompted with a contextual feature query (Fig.  21) which remained constant during the query round.
@@
After completing a round of query response, participants completed post-query questionnaires and inspected the learned models (Fig. 19. Three calibrated feature models were trained while the participant took a 1-2 minute break. The participant then inspected the three learned models and the base feature, responded to the Likert scale for each model, and ranked the four models.
The full experiment script, training presentation, and survey can be found in supplementary material. We provide one survey ordering, where in the other the experiment features are swapped.
Figure 20: Example contextual preference prompt. Short answer descriptions of personal preferences were collected. Despite general priming, participants showed variety in responses.
Figure 21: Example contextual feature query language. Query stayed constant throughout round and was displayed next to query response GUI.
@@
Feature order counterbalancing. We counterbalanced the ordering of the two experimental features: half of the participants saw one feature first, and the other half saw the other feature first. When changing the order of these features, we kept the mapping between the calibrated feature values (trained at 0, 25, 50, and 100 for non-training) and their corresponding numbered relationships (1, 2, 3, and 4 for non-training) fixed. There were 6 possible orderings of the first training feature. For each of these, there were 24 possible combinations of the second and third training features—resulting in too many total permutations to fully counterbalance across all three. To manage this, we used all 6 possible orderings of the first training feature, with each ordering appearing twice (once in each of the two experimental feature orderings). The orderings of the second and third training features were selected to be unique across participants within each ordering condition, but were reused across the two conditions. We aimed for approximate balance by rotating which training queries appeared in each position and ensuring that each participant saw a diverse set of training query orders.
Table 5: Grouped preferences for human distance feature.
@@
1 General preference (n=6) "I would like to make sure that a sharper utensil is further from me as possible. A dull utensil is ok if it is closer."
1 General preference (n=6) "The sharper the utensil, the further it should be from the person."
1 General preference (n=6) "I would like sharper objects to be further away from the person and less sharp objects can be closer."
1 General preference (n=6) "The sharper the utensil the further away"
1 General preference (n=6) "I think I would like sharper utensils to be farther away from the human."
1 General preference (n=6) "Utensil sharpness should correlate with the preferred distance to the person."
2 Unique preference "Keep the knife very far away from human, fork can be at half the arms length. Spatula and spoon could be close but not too close (a hands length away)"
3 Unique preference "The spoon and spatula should be fine wherever. I’d like the fork and knife to be outside a range around the person, after which I don’t really care where they are."
4 Unique preference "There should be a general safety zone around the near vicinity of the person that no utensil should cross (personal space). The sharper the object, the larger that safety zone should be."
5 Unique preference "Distance for the spoon and spatula doesn’t matter. The knife needs to be kept further than the fork but past a certain distance it doesn’t matter how far they are."
6 Specific shared preference (n=2) "generally, the sharper the utensil the further away from the person it should be. Spoons and spatulas are roughly equal in sharpness."
6 Specific shared preference (n=2) "Sharper objects should be further away from the human. The spoon and spatula are equally sharp"
Table 6: Grouped preferences for stove distance feature.
@@ 1 General preference (n=6) "The sharper the utensil the further away"
1 General preference (n=7) "the robot’s hand should be father away from the stove as the heat increases."
1 General preference (n=7) "If the stove is hot, I would like the utensil to be further away. If the stove is not hot, the utensil can be closer."
1 General preference (n=7) "The hotter the stove is the further away the hand should be"
1 General preference (n=7) "The robot hand should be further away from a hotter stove"
1 General preference (n=7) "Increasing stove heat should correlate with the preferred distance from the end effector to the stove."
1 General preference (n=7) "If the stove is on, the utensil should not be directly over it; if the stove is off, this is OK. If the stove is hot, the utensil should not be in relative proximity (tapers off with temperature)."
1 General preference (n=7) "I would want the robot hand to be further away from the stove the hotter the stove is."
2 Unique preference "If the stove heat is at the highest keep the EE furthest away. If the stove is on in general It should be a hands length away. If it’s off the EE can be anywhere"
3 Unique preference "If the stove is on at all, the robot should be at least half the countertop width away."
4 Unique preference "Disregarding what material the utensil is made out of, I would say that the utensil (context) does not really matter. There is a circle around the stove, inside of which I would not like the robot hand. Outside of this circle, I don’t really care where the robot hand is."
5 Unique preference "I don’t want the robot to be close to the stove. Past a certain distance, I don’t care where the robot is. Close is probably where I would expect the robot to feel the heat."
6 Unique preference "I would like the robot’s hand to be further away from the stove the hotter the stove is except when there’s a spatula in hand."

### v3

@@ D.1. Procedures
At the start of each query response round, participants were primed with a general preference, prompting them to teach the effects of a specific contextual element on a specific base feature. Participants were then asked to write down their personal contextual preference (see prompt example for stove feature in Fig. 16). While our method supports accounting for the contextual effects of numerous contextual elements, we used this prompting to steer participants to teach the contextual effects of a single element such that we could visualize changes in the affecting contextual element when they visualized the learned models. The query response graphical user interface (GUI) was then pulled up (Fig. 19), and participants were prompted with a contextual feature query (Fig. 17) which remained constant during the query round.
The same state pairs were shown all participants for training both test features in the discretized Utensil environment (see discretization details required for visualization in App. B.2). Additionally, the laptop location was represented in the state, but when collecting queries, the irrelevant laptop was not visualized in order to prevent cluttering of the visualization.
@@
After completing a round of query response, participants completed post-query questionnaires including the SUS and NASA-TLX. Three calibrated feature models were trained while the participant took a 1-2 minute break. The participant then inspected the three learned models and the base feature (Fig. 20), responded to the Likert scale for each model, and ranked the four models.
@@
Feature order counterbalancing. We accounted for the order of the two experimental features and the order in which trained model variants were presented to users. We counterbalanced the ordering of the two experimental features: half of the participants saw one feature first, and the other half saw the other feature first. When inspecting learned calibrated features, participants were presented with four trained model variants (”Relationships” 1-4) which represented calibrated features trained on 0 (base feature), 25, 50, or all 100 labeled queries (not necessarily in that order). Within each subject, we wanted to ensure the that the mapping between the relationship number and the model variant changed between the three features taught (one in familiarization phase and two in experimental phase). For the first feature trained in the familiarization phase, which had only three associated calibrated models (0, 25, 50), there were 3! = 6 possible orderings of those values. For the second and third experimental features, which had four calibrated models each (0, 25, 50, 100), there were 4! = 24 possible orderings of those values. Fully counterbalancing the order for all three features would therefore have required too many total permutations. To manage this, we used all 6 possible orderings of the first training feature, with each ordering appearing twice (once in each of the two experimental feature orderings). The orderings of the second and third training features were selected to be unique across participants within each ordering condition, but were reused across the two conditions. We aimed for approximate balance by rotating which training queries appeared in each position and ensuring that each participant saw a diverse set of training query orders.
Screenshot of prompt for user description of personal contextual preference shown to participants.
Figure 16. Example contextual preference prompt. Short answer descriptions of personal preferences were collected. Despite general priming, participants showed variety in responses.
Screenshot of prompt for user description of personal contextual preference shown to participants.
Example contextual feature query language. Query stayed constant throughout round and was displayed next to query response GUI.
Figure 17. Example contextual feature query language. Query stayed constant throughout round and was displayed next to query response GUI.
Screenshot of contextual feature query language shown to participants.
Examples of two different user contextual preferences for the stove feature, depicted by the user’s written preference description the learned calibrated feature. Learned calibrated features are evaluated at all states in a point cloud and shown at different levels of stove heat.
Figure 18. Differences in user preferences captured in learned calibrated features.
Examples of two different user contextual preferences for the stove feature, depicted by the user’s written preference description the learned calibrated feature. Learned calibrated features are evaluated at all states in a point cloud and shown at different levels of stove heat.
D.2. Graphical user interfaces
Participants interacted with two GUIs in order to both respond to contextual feature queries (Fig. 19), and to interactively inspect learned calibrated features (Fig. 20). When responding to queries, participants were able to view the states from different angles, and in response to the visualized contextual feature query select between ”Select State 1”, ”Select State 2”, or ”States are equal”. Users could also switch back and forth between the two compared states, and the current state was visualized in order to help users stay aware. Additionally, environment objects, like utensils, were colored differently to help make relevant context immediately salient and speed up query response.
Learned models were visualized as described in Fig. 2; however, 5k states were used as opposed to 8k for speed of evaluation and visualization. Participants could switch between the four models, and visualize a model in varying contexts by dragging a sliding bar which snapped the visualized context to the nearest of four spread out discrete contexts (and changed the model visualization accordingly). Participants were instructed on helpful ways to investigate and compare models, such as inspecting a single model across contexts, and switching between all models to see how they compare in a single context. Participants provided no input in this GUI, instead using it to respond to a survey collecting the Likert scale evaluation for each model and model rankings.
D.3. Preference analysis
Each participant provided a short free-response description of their contextual preference prior to responding to contextual feature queries (see prompt in Fig. 16). Therefore, we collected 12 short free-response answers regarding contextual preferences for each of the two test features. A single reviewer manually reviewed the responses and conservatively grouped them based on the level of specificity provided (Tables 34). Responses describing contextual preferences in general, high-level terms without specific details were grouped together, while responses including concrete details were grouped separately unless the details matched exactly. This inductive grouping was intended to get a preliminary understanding of preference diversity. We found 6 users shared a general preference for the human feature, with 2 users sharing a similar variant and 4 users holding unique variants. For the stove feature, we found 7 users shared a general preference, and the remaining 5 users described individual variations.
In Fig. 18 we give another example similar to Fig. 6, where we demonstrate how learned calibrated features differ for users with different preferences. The main difference in this example is the context in which the stove is off–while for one user the feature is equally respected in all states when the stove is off, the other user still desires the robot to stay away from directly above the stove.
Screenshot of graphical user interface. Shows manipulation environment robot and control panel.
Figure 19. Query response GUI showing one state in pair. Participants used buttons on right panel to switch between visualizing two states and respond to queries. Left hand panel can be ignored.
Screenshot of graphical user interface. Shows manipulation environment robot and control panel.
Screenshot of graphical user interface. Shows point cloud of learned calibrated feature overlaid on manipulation environment and control panel.
Figure 20. Model visualization GUI. Calibrated feature for stove feature trained with 50 queries visualized in high heat context. Participants used buttons on right panel to select different models to visualize and sliding bar to move between four discrete contexts. Left hand panel can be ignored.
Screenshot of graphical user interface. Shows point cloud of learned calibrated feature overlaid on manipulation environment and control panel.
Table 3. Grouped preferences for human feature.
@@ D.3. Preference analysis
1 General preference (n=6) ”I would like to make sure that a sharper utensil is further from me as possible. A dull utensil is ok if it is closer.”
1 General preference (n=6) ”The sharper the utensil, the further it should be from the person.”
1 General preference (n=6) ”I would like sharper objects to be further away from the person and less sharp objects can be closer.”
1 General preference (n=6) ”The sharper the utensil the further away”
1 General preference (n=6) ”I think I would like sharper utensils to be farther away from the human.”
1 General preference (n=6) ”Utensil sharpness should correlate with the preferred distance to the person.”
2 Unique preference ”Keep the knife very far away from human, fork can be at half the arms length. Spatula and spoon could be close but not too close (a hands length away)”
3 Unique preference ”The spoon and spatula should be fine wherever. I’d like the fork and knife to be outside a range around the person, after which I don’t really care where they are.”
4 Unique preference ”There should be a general safety zone around the near vicinity of the person that no utensil should cross (personal space). The sharper the object, the larger that safety zone should be.”
5 Unique preference ”Distance for the spoon and spatula doesn’t matter. The knife needs to be kept further than the fork but past a certain distance it doesn’t matter how far they are.”
6 Specific shared preference (n=2) ”generally, the sharper the utensil the further away from the person it should be. Spoons and spatulas are roughly equal in sharpness.”
6 Specific shared preference (n=2) ”Sharper objects should be further away from the human. The spoon and spatula are equally sharp”
Table 4. Grouped preferences for stove feature.
@@ 1 General preference (n=6) ”The sharper the utensil the further away”
1 General preference (n=7) ”the robot’s hand should be father away from the stove as the heat increases.”
1 General preference (n=7) ”If the stove is hot, I would like the utensil to be further away. If the stove is not hot, the utensil can be closer.”
1 General preference (n=7) ”The hotter the stove is the further away the hand should be”
1 General preference (n=7) ”The robot hand should be further away from a hotter stove”
1 General preference (n=7) ”Increasing stove heat should correlate with the preferred distance from the end effector to the stove.”
1 General preference (n=7) ”If the stove is on, the utensil should not be directly over it; if the stove is off, this is OK. If the stove is hot, the utensil should not be in relative proximity (tapers off with temperature).”
1 General preference (n=7) ”I would want the robot hand to be further away from the stove the hotter the stove is.”
2 Unique preference ”If the stove heat is at the highest keep the EE furthest away. If the stove is on in general It should be a hands length away. If it’s off the EE can be anywhere”
3 Unique preference ”If the stove is on at all, the robot should be at least half the countertop width away.”
4 Unique preference ”Disregarding what material the utensil is made out of, I would say that the utensil (context) does not really matter. There is a circle around the stove, inside of which I would not like the robot hand. Outside of this circle, I don’t really care where the robot hand is.”
5 Unique preference ”I don’t want the robot to be close to the stove. Past a certain distance, I don’t care where the robot is. Close is probably where I would expect the robot to feel the heat.”
6 Unique preference ”I would like the robot’s hand to be further away from the stove the hotter the stove is except when there’s a spatula in hand.”